### v1

# Efficiently Ranking Software Variants with Minimal Benchmarks

Abstract.
Benchmarking is a common practice in software engineering to assess the qualities and performance of software variants, coming from multiple competing systems or from configurations of the same system. Benchmarks are used notably to compare and understand variant performance, fine-tune software, detect regressions, or design new software systems. The execution of benchmarks to get a complete picture of software variants is highly costly in terms of computational resources and time. In this paper, we propose a novel approach for reducing benchmarks while maintaining stable rankings, using test suite optimization techniques. That is, we remove instances from the benchmarks while trying to keep the same rankings of the variants on all tests. Our method, BISection Sampling, BISS, strategically retains the most critical tests and applies a novel divide-and-conquer approach to efficiently sample among relevant remaining tests. We experiment with datasets and use cases from LLM leaderboards, SAT competitions, and configurable systems for performance modeling. Our results show that our method outperforms baselines even when operating on a subset of variants. Using BISS, we reduce the computational cost of the benchmarks on average to 44% and on more than half the benchmarks by up to 99% without loss in ranking stability.
benchmark,minimization, black-box,cost reduction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering Software performanceccs: Software and its engineering Software configuration management and version control systemsccs: Software and its engineering Software testing and debugging
1. Introduction
Figure 1. High Level Overview of Ranked Test Suite Minimization (RTSM) which aims at minimizing benchmarks while keeping the ranking of the variants corrects enabling reliable, fast and cheap evaluation of new variants.
[…]
The figure shows a set of variants and 100 tests being reduced by solving the RTSM problem to 45 tests while producing the same overall ranking of the variants.
Benchmarking is a fundamental practice in software engineering that focuses on evaluating the qualities and performance of software systems. It is also essential in computer science for tasks such as optimization of algorithms or evaluation of system performance (Bartz-Beielstein et al., 2020; Olson et al., 2017; Kerschke et al., 2019; Malan and Engelbrecht, 2013; Muñoz Acosta et al., 2015; Moreau et al., 2022). In computational science, scientists use benchmarking to explore, compare and validate the performance of complex variants of simulations and models, facilitating advances in research and development (Donoho, 2024; Guyon and Viegas, 2020). Benchmarks can be seen as a special kind of test that is used to measure software performance such as execution time, energy consumption, or accuracy. Unlike unit or integration tests that return binary or ternary test results (i.e., Pass/Fail/inconclusive), benchmarking returns numerical evaluations, which makes them more challenging to work with. This is why software system benchmarking usually involves comparisons with other software variants. Hence, benchmarks are executed on different variants, whether they originate from different competing, alternative systems, or various configurations of the same system.
These benchmarks, workloads, test sets, performance tests, inputs, or problems typically consist of a set of test cases to measure and compare the software variants’ performance. The ranking of software variants based on benchmark results provides an objective quantitative measure of their relative performance. This ranking not only highlights discrepancies in variant performance, but also offers a concrete basis for comparing the diversity of software solutions across different problem domains. In addition, rankings can reveal patterns and trends in performance across various types of inputs or problem characteristics, providing valuable, data-driven insights for software optimization and development strategies.
Benchmarks and performance comparisons are crucial for understanding the performance characteristics of software variants, configuring and fine-tuning software, detecting regressions, and guiding the design of new systems. However, executing comprehensive benchmarks to achieve a complete understanding of software variants can be extremely costly in terms of computational resources and time. For instance, LLM leaderboards (Liu et al., 2023a) or SAT competitions (Froleyks et al., 2021; Heule et al., 2019) are two well-known examples of benchmarking competitions that require extensive computational resources to evaluate the performance of variants. The need for more efficient methods to maintain accurate rankings of software variants while minimizing the computational load is evident in several concrete scenarios:
  • Comparison: In competitive settings or for the sake of improvements, numerous software variants must be evaluated and ranked. For instance, leaderboards are becoming more and more available to provide engineers with a variety of benchmarks and the associated performance of the variants. Full benchmarks are often prohibitively expensive, especially with many entries. A minimized benchmark significantly reduces computational costs while still accurately ranking top performers and revealing their strengths and weaknesses.
  • Regression testing: Executing reduced benchmarks on each commit or release allows for frequent performance checks, enabling quick detection of regressions’ rankings, for example, in continuous integration. This approach facilitates rapid iterations and more thorough exploration of alternatives in both software engineering and scientific computing, with better-informed decisions and continuous performance assessment;
  • Performance modeling, extrapolation, and configuration tuning: Engineers often use mathematical models and simulations to predict software performance under various conditions or to find optimal configurations (Balsamo et al., 2004). By selecting a subset of representative benchmarks, they can efficiently train models that capture essential performance distributions. When benchmarking is training (Bartz-Beielstein et al., 2020), this process reduces the need for exhaustive execution, allowing engineers to efficiently identify performance bottlenecks and optimize system configurations based on workload sensitivity (e.g., as recently shown in (Mühlbauer et al., 2023; Lesoil et al., 2023)).
  • Benchmark design and elaboration: In case benchmarks are repetitive or identical, a strategy to design better benchmarks can be to focus on the most informative tests and extend them;
  • Benchmarking a new system: When new systems (variants) are considered, reduced benchmarks can be leveraged at a better cost to get a preliminary performance comparison with competitors.
In short, our goal is to equip software engineers, computational scientists, and researchers with an objective method to reduce the computational cost of benchmarks while still effectively assessing competing software-based systems (see Fig. 1). It can boost innovation and productivity by enabling more efficient evaluation or simply reducing the cost of benchmarking.
However, all of these scenarios also highlight the potential dangers of reduced benchmarks that do not fully represent the entire set of benchmarks. To our knowledge, no previous work has addressed the problem of minimizing benchmarks while preserving the ranking of multiple software variants (see also more details in section 2). Previous research has extensively explored benchmarking and testing (Yoo and Harman, 2012), focusing on optimization, prioritization, and reduction of the test suite, but primarily for single systems, functional tests, and not for competing software variants. The works on software product lines (SPL) focus on the analysis of multiple products (Lima et al., 2020; Al-Hajjaji et al., 2017; Elbaum et al., 2014; Ensan et al., 2011; Kumar and Rajkumar, 2016; Lachmann et al., 2017, 2015; Lity et al., 2019; Lochau et al., 2012; Marijan et al., 2019, 2013; Marijan and Liaaen, 2017; Marijan et al., 2017; Prado Lima and Vergilio, 2020; Runeson and Engström, 2012; Wang et al., 2015, 2016; Xu et al., 2013), but assume a single code base and seek to minimize the number of test cases while usually maximizing the functional coverage of features or code (not performance).
Our optimization goal thus differs and consists in preserving the performance ranking of potentially independent software variants. In this paper, we propose a novel method to reduce the computational cost of benchmarks while still effectively ranking software variants using test suite optimization techniques. Our method, called BISS for BISection Sampling, is based on the fact that some tests cannot be removed and form the essential core of the benchmark. Furthermore, once some tests are removed, other tests become crucial as well. This enables us to use a bisection sampling routine to sample among relevant tests cleverly. We couple it with a divide-and-conquer algorithm to reduce the variance and speed up the search, and with an automatic restart (iterative solving) as long as we find improvements. Our contributions include:
  • The introduction of a new problem in Software Testing research, that we frame as the ranked test set minimization problem (RTSM);
  • A novel divide-and-conquer method exploiting a prediction model and parameterized search to practically solve the RTSM problem and effectively prune test suites;
  • Experimental results showcasing, on 50 scenarios and more than 30 different systems, that our approach outperforms multiple baselines and its impacts, enabling the removal of 20% up to 99% of tests depending on the accuracy requested;
  • An artifact with all the datasets and benchmarks needed to reproduce the results.
The remainder of this paper is organized as follows: section 2 discusses related work further. In section 3 we introduce the problem. In section 4 we first describe our method and then in section 5 propose baselines to compare to, including implementation details. We provide experimental results in section 6. Then we focus on selected case studies in  section 7. We discuss threats to validity in section 8. We discuss results and provide recommendations for developers and benchmark designers and users in section 9. Finally, we draw general conclusions and discuss future work in section 10.
2. Related work
Numerous work on regression testing and test suite minimization, selection, and prioritization has been published (Yoo and Harman, 2012). The principle is to improve the test suite of a single software system, typically to reduce computational cost as test suites tend to grow in size. Test suite minimization seeks to prune test suites from redundant test cases to reduce the number of tests to run. We pursue a related, yet different, goal with two key differences: we are interested in performance of software and we consider multiple software variants.
Research in software performance engineering explores benchmarking at different levels (e.g., system-level tests and method/statement-level tests) and with different approaches (e.g., blackbox vs whitebox) (Laaber et al., 2020; Weyuker and Vokolos, 2000; Jiang and Hassan, 2015; Menascé, 2002; Blot and Petke, 2025; Nguyen et al., 2014; AlGhamdi et al., 2020). For example, there are techniques for microbenchmarking that measure the performance of e.g., single functions (Laaber et al., 2020; Maricq et al., 2018; Laaber et al., 2020; Costa et al., 2019). Ding et al. (Ding et al., 2020) studied how unit tests can be used to assess performance properties. All of these works do not aim to preserve the performance ranking of multiple variants.
In close proximity to our work are contributions related to software product lines (SPLs) and configurable systems (Thüm et al., 2014; Rhein et al., 2018; Lima et al., 2020; Al-Hajjaji et al., 2017; Elbaum et al., 2014; Ensan et al., 2011; Kumar and Rajkumar, 2016; Lachmann et al., 2017, 2015; Lity et al., 2019; Lochau et al., 2012; Marijan et al., 2019, 2013; Marijan and Liaaen, 2017; Marijan et al., 2017; Prado Lima and Vergilio, 2020; Runeson and Engström, 2012; Wang et al., 2015, 2016; Xu et al., 2013; Chen et al., 2019). In this line of research, the goal is to consider a set of related products or configurations coming from a common code base (an SPL or a configurable system) and to optimize testing activities for these products. Existing SPL works explore the three traditional regression testing techniques for single software systems (Yoo and Harman, 2012), namely minimization (Wang et al., 2015), selection (Lity et al., 2019; Lochau et al., 2012; Wang et al., 2016; Xu et al., 2013), and prioritization (Al-Hajjaji et al., 2017; Lachmann et al., 2017, 2015) of test cases. Wang et al. propose to optimize test suites considering effectiveness measures such as test minimization percentage, pairwise coverage, fault detection capability (Wang et al., 2015) but these works ignore software performance properties and ranking information. Mendonça et al. introduce FeaTestSelPrio a feature-oriented test case selection and prioritization approach that improves failure detection and reduces test execution time by linking test cases to feature implementations and prioritizing based on feature coverage (Mendonça et al., 2024). Our work differs from several perspectives: (1) we target variants from multiple code bases; (2) we leave apart variant features: such features can be unavailable in our context, where we only have access to the software variants themselves; (3) performance calls to consider quantitative properties of software, whereas test suites are typically qualitative (e.g., Fail/Pass or Inconclusive); (4) our optimization goal is to preserve the (performance) ranking of software variants, while prior works consider failures or feature/code coverage.
Instead of reducing benchmarks, a more common problem is the selection of relevant software configurations for a given problem. Several configuration sampling strategies have been proposed in the literature on SPLs and configurable systems, either to find and cover as many faults as possible (Han and Yu, 2016; Thüm et al., 2014), or to measure and learn the performance of the configuration space (Alves Pereira et al., 2021, 2020; Kaltenecker et al., 2020). Their focus is on the selection of configurations (variants) to test, rather than on the selection of test cases to run. Recent empirical findings from Mühlbauer et al. (Mühlbauer et al., 2023) and Lesoil et al. (Lesoil et al., 2023) show that program inputs (or workloads), typically part of a benchmark, and configuration options can result in substantial performance variations. The sheer volume of configurations and benchmark elements creates high computational costs, making it impractical to learn a performance model for each test case. These performance variations can alter the ranking of software variants, highlighting the need for careful benchmark selection to ensure accurate assessments. Nair et al. (Nair et al., 2017) tackle configuration optimization through rank-based selection assuming a given workload and relying on potentially imprecise performance models. In contrast, our approach focuses on reducing the benchmark suite itself to preserve the ranking across software variants, rather than identifying optimal configurations. Existing work, such as Nair et al., is not applicable to benchmark reduction, as it focuses on optimizing configurations within a single workload and does not address the challenge of selecting benchmarks to maintain ranking consistency across multiple variants.
Multimorphic testing evaluates the coverage of a test suite with respect to a quantitative property by generating multiple software variants to reveal performance differences (Temple et al., 2018). Assesses performance test suite effectiveness by identifying tests that can ”kill” underperforming configurations and calculating a dispersion score to measure coverage across variants. While we operate in a similar context, our focus differs fundamentally: instead of the dispersion score, we aim to ensure that the minimized benchmark set preserves the correct performance ranking among competing variants, which is essential for various scenarios described above.
The automated algorithm selection problem is subject to intensive research (Kerschke et al., 2019; Hutter et al., 2011, 2014; Xu et al., 2008; Thornton et al., 2013) with many applications in artificial intelligence (such as SAT, CSP, QBF, ASP, or scheduling (Kerschke et al., 2019)). Given a computational problem, a set of algorithms, and a specific problem instance to be solved, the challenge is to determine which of the algorithms can be selected to perform best on that instance. For example, SATZilla uses machine learning to select the most effective algorithm from a portfolio of SAT solvers for a given SAT formula (Xu et al., 2008). In addition, determining which of two algorithms performs best on a benchmark, without running the whole benchmark, has been discussed in (Matricon et al., 2021). However, the focus is on two variants and the cost of running variants, not on benchmarks. A critical factor in automated algorithm selection is the set of problem instances (benchmarks) used for e.g., training models to facilitate optimal algorithm selection. Our work contributes to this area by providing a software engineering methodology that can be applied to various AI problems.
3. Problem Statement: Ranked Test Suite Minimization (RTSM)
In software testing, test suite minimization focuses on reducing the number of tests needed while still ensuring that all requirements are met. Traditionally, these functional tests check Boolean requirements, where a test either passes or fails, similarly to solving set-cover instances. However, many scenarios involve performance-related tests that have continuous and quantitative answers.
Definition 3.0.
Given a set of configuration variants $\mathcal{A}$, a performance value is a function $t:\mathcal{A}\to\mathbb{R}$ that associates a performance value (e.g. estimation, measurement, execution time, memory consumption, etc.) with each variant of a configurable system.
Individual performance values are usually difficult to get and verify. For example, checking that a system runs exactly in 2.42 seconds, as it has been registered from previous executions or predicted by an ML model, is very unlikely. Comparison of this with other observations coming from other variants’ performance estimations is usually needed. Ranking performance estimations are a way to evaluate how a system operates and is widely used (Bartz-Beielstein et al., 2020; Lesoil et al., 2023).
Given a set of configuration variants, we can typically estimate their performance on a set of test inputs, sum their performance over these inputs, and rank each variant accordingly. Intuitively, given a full benchmark obtained from each variant’s performance estimation and its ranking of the variants, our goal is to compress the benchmark while preserving the ranking.
Definition 3.0.
The rank of a variant $v\in\mathcal{A}$ for a test suite $\mathcal{T}$ is the rank $r$ of its performance estimation $perf(v,\mathcal{T})=\sum_{t\in\mathcal{T}}t(v)$, precisely:
$r=Index(\operatorname*{\arg\!\max}_{v\in\mathcal{A}}(perf(v,\mathcal{T}))$
The variant with the maximal performance estimation is ranked $1$ and so on. Note that this ranking computes a total ordering over the variants in $\mathcal{A}$.
Definition 3.0.
Given a set of variants $\mathcal{A}$ and a test suite $\mathcal{T}$, the Ranking Test Suite Minimization problem (RTSM) aims to find a smallest $\mathcal{T}^{\prime}\subset\mathcal{T}$ and $W$ such that the ranking of variants $\mathcal{A}$ for $\mathcal{T}^{\prime}$ is preserved (i.e., the same as the ranking of variants $\mathcal{A}$ on $\mathcal{T}$) where $W$ gives a weight to each test in $\mathcal{T}^{\prime}$ for ranking. Note that several subsets $\mathcal{T}^{\prime}$ can be solutions of RTSM.
Looking at the left part of the Fig. 1, we have an initial set of variants along with the tests set and the performance estimation of each pair variant test. Their ranking is determined by classifying the variants in ascending order with respect to the summed performance estimates. A solution to the RTSM problem is given on the right part of the figure. Keeping only 45 tests as a subset of tests, the ranking of variants is preserved even though the total performance estimation has changed.
Note that test suite minimization is an instance of RTSM when one just ignores the ranking. As test suite minimization is NP-hard (Gotlieb and Marijan, 2014), we can deduce from the previous remark that RTSM is also NP-hard. In practice, for large instances, we will accept near-optimal solutions obtained by using heuristics, anytime algorithms, or probabilistic approaches.
Definition 3.0.
Given a set of variants $\mathcal{A}$ and a test suite $\mathcal{T}$ with a cost associated with each test, the Weighted Ranking Test Suite Minimization problem (WRTSM) aims to find $\mathcal{T}^{\prime}\subset\mathcal{T}$ and $W$ such that the ranking of variants $\mathcal{A}$ for $\mathcal{T}^{\prime}$ is preserved (i.e., the same as the ranking of variants $\mathcal{A}$ on $\mathcal{T}$) and that minimizes $cost(\mathcal{T}^{\prime})=\sum_{t\in\mathcal{T}^{\prime}}cost(t)$ where $W$ gives a weight to each test in $\mathcal{T}^{\prime}$ for ranking.
Error Metric
Unfortunately, there is no ground truth; therefore, we will compare the rankings obtained using the solution to the original solution. There are multiple metrics to compare the rankings, the two most commonly used ones are Kendall’s (Kendall, 1938) and Spearman’s (Spearman, 1904) rank correlation coefficient. Both give a score between -1, the two vectors offer opposite rankings, and 1, the two vectors offer exactly the same rankings. We chose to use Kendall’s, which is directly proportional to the number of variants correctly ranked. Choosing Spearman or Kendall does not matter when the aim is a score of 1, that is, the same ranking, because they both coincide for this data point. Both tests are non-parametric hypothesis tests on the relations between the two rankings. Kendall can be easily computed, assume that a pair of variants are concordant if they are ranked in the same order in both rankings and discordant otherwise, then we have:
$\text{Kendall's }\tau=\frac{\#|concordant\ pairs|-\#|discordant\ pairs|}{\#|pairs|}$
Special cases occur when there are ties, but in our scenarios, ties never occur.
Multi-metric
The current formulations in Definition 3.2 and Definition 3.3 use a single performance metric, but can be easily adapted to multiple metrics. When multiple metrics are present, compute the Kendall for each metric independently and take the worst one. For example, if we wanted a Kendall of .99 in the multi-metric case, then on all metrics the Kendall must at least be .99.
4. Benchmark minimization with BISS
Novelty
As explained in section 2, most existing methods for minimizing the test suite require prior information such as test results, historical data, etc. However, the scenarios for RTSM are very diverse, requiring handcrafted features by experts for their domain if we want to use a white-box approach, which are often not available or are too costly to acquire. Instead, we choose to look at black-box approaches. We have no features about the variants; we have just an identifier to be able to identify them, and this is the same for the tests. So, as input, we only have the performance matrix, that is, a matrix that contains the performance of each variant on each test. Then techniques such as principal component analysis (PCA) or racing techniques (López-Ibáñez et al., 2016) should solve this problem. We will see that we tested these techniques and they do not solve our problem. Indeed, these classic techniques can work to some extent, but we are interested in objectives that are a loose proxy of our interest: the ranking of variants. Whereas most of these techniques are interested in more fine-grained information: the magnitude of the differences, they care if you are twice as fast or ten times as fast, whereas in RTSM it does not matter.
Intuition
The core idea is that we do not care about the individual performances of the tests; we only care about the relative performance of a variant to another across tests.
Definition 3.2 gives us the possibility of assigning weights to $W$: Given a subset of the test set, we can use a linear regression to optimize $W$ to obtain the best coefficients. Let $P\in\mathbb{R}^{|Variants\times|\mathcal{T}|}$ be the performance matrix of the variants in the test, then, given $\mathcal{T}$’ of $\mathcal{T}$, we can optimize:
$\forall i\in\text{Variants, }P_{i\mathcal{T}^{\prime}}W^{T}=\sum_{j=0}^{|\mathcal{T}|}P_{ij}$
This is a linear regression problem where $W$ is a vector that assigns a weight to each test in $\mathcal{T}^{\prime}$, which is the set of parameters that we try to find for each variant, given their performance in the variants, to predict their performance on all tests. Our goal is therefore to sample relevant subsets $\mathcal{T}^{\prime}$. Formally, we assume that we have the procedure $solves_{T}(T^{\prime})$ which solves the linear regression and checks if the Kendall target is reached with $T^{\prime}$ as a solution to the problem with the complete test set being $T$, if it is, then the procedure returns truetrue otherwise, it returns falsefalse.
As we have explained earlier, the problem is NP-hard, and covering all possible solutions is intractable for small test suites; for a test suite of size 30, there are more than a billion possible solutions. The process is a bit akin to dimensionality reduction; each test acts as a dimension, and we want to reduce the dimension while keeping the ranks intact. Taking inspiration from the Johnson–Lindenstrauss lemma, which states that a random projection in the ”low” dimension has a high probability of keeping the original distances between the data points in the original high dimension. The dimension that the lemma gives us is dependent logarithmically on the original dimension and grows inversely proportional to the square of the minimum distance between two data points. In other words, the closer two variants have similar performance, the more likely it is that we are going to need more tests to rank them correctly. Therefore, our objective is to sample subsets of tests intelligently that are likely to result in a solution.
4.1. Bisection Sampling
Obviously, we would like to sample more cleverly than pure random. For that, we need to improve on two points: reducing the sampling variance and eliminating tests faster than one at a time.
Variance reduction
The idea is to draw on the following fact: If we add more tests to a solution to the RTSM problem, then it is also a solution. Then, conversely, if we have a set of tests that is not a solution, none of the subsets is a solution. Given a solution $T$ for each test $t\in T$ we can check if $solves_{T}(T\setminus\{t\})$ is truetrue, that is, if with all the information except for test $t$ we have a solution. If $solves_{T}(T\setminus\{t\})$ is falsefalse then it must be that the information of test $t$ is necessary if we want to reduce $T$ further; therefore, it is pointless to try to remove it. Iterating this process over all tests of $T$ partitions the tests into two sets: a set of tests $N$ that are necessary, and a set of tests $P$ that can potentially be removed. This process is described in Algorithm 1. Notice that it is quite costly because we need to do this for every test that can be potentially removed, but in practice, we hope that it reduces the number of elements we have to sample from.
Algorithm 1 FindNecessary($P$, NN, $\mathcal{T}$): find the tests among a set of tests $P$ that are necessary in order to reduce $P\cup N$ for the test suite $\mathcal{T}$
1:for $t\in P$ do
2:  $C\leftarrow(N\cup P)\setminus\{t\}$
3:  if not $solves_{\mathcal{T}}(C)$ then
4:    $N\leftarrow N\cup\{t\}$
5:    $P\leftarrow P\setminus\{t\}$
6:  end if
7:end for
8:return $N,P$
Fast Elimination
The idea is to use a bisection algorithm as in (Bentley et al., 2019), for example, where a bisection algorithm is used to find the relevant code locations. Bisection algorithms are similar to binary search; the general idea is to use a divide-and-conquer approach where the search space is split, then tested, and based on the results of the test, it is reduced, and then the process repeats until one element is found. Here we are looking for one set among a set of test sets. Here we have $P$, the set of tests from which we want to sample, $N$, the set of tests that are necessary, and $\mathcal{T}$ the test suite considered, but the set of possible solutions is too large, making the bisection algorithm impractical. Instead, the idea is to randomly divide $P$ into two sets $A$ and $B$ of the same size. Then we can do the following: check if $A$ or $B$ is a solution with $solves_{\mathcal{T}}(A\cup N)$, if that is the case, we divide the size of $P$ by a factor of two, and we can repeat the process. If none of them are solutions, then we make the following assumption that either $A$ or $B$ is necessary; therefore, we compute the solution where $A$ is necessary and the solution where $B$ is necessary and return the best. In each step, we divide the size of $P$ by two, which allows much faster progress.
Algorithm
This process is formalized in Algorithm 2. The first step is to use Algorithm 1 to reduce the variance, and then we can proceed with the sampling.
Algorithm 2 BS($P$, NN, $\mathcal{T}$): sample a solution to the RTSM problem for $\mathcal{T}$
[…]
2:if $P$ is empty then
3:  return NN
4:end if
[…]
7:if $solves_{\mathcal{T}}(A\cup N)$ then
8:  return $BS(A,N,\mathcal{T})$
9:else if $solves_{\mathcal{T}}(B\cup N)$ then
10:  return $BS(B,N,\mathcal{T})$
11:else
12:  $S_{A}\leftarrow BS(B,A\cup N,\mathcal{T})$
13:  $S_{B}\leftarrow BS(A,B\cup N,\mathcal{T})$
14:  return best of $S_{A}$ or $S_{B}$
15:end if
4.2. Divide and Conquer
In reality, $solves_{K\cup P}(K)$ does a linear regression, which complexity is O($|\mathcal{A}||K|^{2}|P|)$ where $\mathcal{A}$ is the set of variants. Therefore, testing for a large test set $K$ becomes very costly, as in $FindNecessary$, where we remove a single test. Thus, a more scalable approach needs to be used.
[…]
Figure 2. High-level overview of our divide and conquer with an iterative strategy used in our BISS algorithm
[…]
The figure illustrates a classic divide and conquer approach on random subsets of the input test set except that if the solution subset is smaller than the original the algorithm is restarted with the solution.
The idea is to use a divide-and-conquer approach and to use bisection sampling as a subroutine. Fig. 2 illustrates at a high level our divide-and-conquer approach with our iterative strategy. In the first step, we split $\mathcal{T}$ into multiple smaller sets and solve these smaller instances of the RTSM problem. From then on, at each iteration, we will merge two solutions of the previous step. We do this simply by considering the instance of the RTSM problems, where we keep only the tests in either solution. This is a new instance of the RTSM problem that we solve. We keep iterating until there is only one instance left. In the worst case, in all of the sub-instances, we could not find better solutions than including every test of the sub-instance; therefore, the last step is similar to not doing any divide and conquer, but in all other cases, we potentially solve smaller problems, enabling faster solving. The initial split thus has a large variance and, therefore, can have a large impact on the performance of this approach.
The process is described in detail in Algorithm 3. But before diving into the explanations of the algorithm, let us convince ourselves that this process gives a valid solution. Given two solutions $A$ and $B$ for the RTSM instances $T_{A}$ and $T_{B}$, is $S$ a solution to the RTSM instance $A\cup B$ also a solution to the entire RTSM instance $T_{A}\cup T_{B}$? The answer is no, the reason being that when we build the predictions, we build them on the exact true data, so when we consider only $A\cup B$ we may be able to remove the tests because we can predict them accurately up to their ranks but not accurately enough for use in the prediction of other tests. To do this correctly at minimal cost, when a solution $S$ is found to the RTSM instance $A\cup B$ we check if it is also a solution to the entire RTSM instance $T_{A}\cup T_{B}$, if it is we accept it and continue otherwise we try again a certain number of times until we arbitrarily set $S$ to $A\cup B$.
Algorithm 3 DC($\mathcal{T}$, $n$, SampleSample): from a set of tests $\mathcal{T}$ by splitting it into $n$ parts and using the SampleSample to sample solutions, find a solution by using divide and conquer
[…]
3:while $Q$ is not empty do
4:  $(A,T_{A}),\ (B,T_{B})\leftarrowfirst$ two elements of $Q$
5:  $S\leftarrow Sample(A\cup B,\emptyset,A\cup B)$
6:  if not $solves_{T_{A}\cup T_{B}}(S)$ then $\triangleright$ check if $S$ is NOT a valid merge
7:    $S\leftarrow A\cup B$ $\triangleright$ discard the intermediate solutions
8:  end if
9:  $Q.append((S,T_{A}\cup T_{B}))$
10:end while
11:return $pop(Q)$
Finally, we combine our divide-and-conquer with iterative solving, since our method is based on random sampling. Iterative solving solves the RTSM problem from an already existing solution and follows the idea that if a smaller cost test set is found, then we may try one more time to find another improvement until no improvement is found.
5. Other Approaches
We describe a few methods inspired by related work and intuition that we consider as baselines. They can also be used as divide-and-conquer subroutines, including iterative solving, algorithms, and will try solutions and test them with solvessolves.
Our method is called BISS. All methods, unless specified, use divide-and-conquer and iterative solving.
Random Search RANDOM
Random search is a simple baseline that can be easily adapted to the RTSM problem. The random search involves randomly sampling a set of tests and then checking if the set forms a solution that can be returned at any time. Rather than sampling among all possible candidates, the process can be adapted to be iterative. Initially, there are $n$ tests in the suite, then we randomly search for a set of tests of size n1n-1.
Greedy Algorithm GREEDY
Another simple black-box baseline in a greedy search approach. At each step, we can try to remove the test that provides the least information; since all tests have the same weight, we chose to remove the test with the lowest metric variance because, on average, it should contribute the least to the ranking. Note that we also tried with the lowest mean metric, and the results were approximately the same, so we decided not to report this variant for brevity reasons.
Mixed Integer Linear Programming MILP
The WRTSM problem can be transformed into an MILP problem, except that since our decision variables are the presence or absence of a test in the solution, this disables the ability to also optimize $W$, which weights the tests in the subsets. Let us use the notations introduced in Definition 3.3: $\mathcal{A}$ is the set of variants, $\mathcal{T}$ is the set of tests, $p_{k,i}$ if the performance of variant $k$ in the test $i$, $cost(t_{i})$ is the cost of the test $t_{i}$ and the weights of the solution $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$$\in\mathbb{R}^{+}$ for the test $t_{i}$. Let us explain our encoding; the presence or absence of the test $t_{i}\in\mathcal{T}$ in the solution is encoded as ${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\in\{0,1\}$. Our constraints are encoded as follows:
$\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{k,i}$ $\leq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{l,i}$ if variant $k$ is ranked lower than variant $l$
$\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{k,i}$ $\geq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{l,i}$ if variant $k$ is ranked higher than variant $l$
which translates to the ranking should be the same. Thus, our objective is to minimize:
$\min\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times cost(t_{i})$
However, this formulation does not take into account the weights of the tests in the solution. They can be included in the following way into our constraints for variant $k$ and variant $l$:
$\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}{\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}\times p_{k,i}\leq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}{\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}\times p_{l,i}$ if variant $k$ is ranked lower than variant $l$
$\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}{\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}\times p_{k,i}\geq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}{\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}\times p_{l,i}$ if variant $k$ is ranked higher than variant $l$
Notice that this is no longer a linear program, since both the $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} and the $\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}$} are decision variables. Therefore, we use the previous linear version even though it solves a simpler version of the problem, but with less agency; therefore, the solutions of the MILP approaches are not optimal for the problem where the $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} can be set, although they are optimal for the ${\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}=1$ cases. Observe that this formulation also only enables us to consider a Kendall of exactly 1, if we wanted to consider a Kendall of $T$, then we would use the following formulation for variant $k$ and variant $l$:
${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{k,i}\leq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{l,i}$ if variant $k$ is ranked lower than variant $l$
${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{k,i}\geq\sum_{i}{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\times p_{l,i}$ if variant $k$ is ranked
with ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}\in\{-1,1\}$, they are also decision variables and the additional constraint that:
$\frac{1}{2|\mathcal{A}|^{2}}\sum_{k,l\in\mathcal{A}}{\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}\geq T$
since ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=1$ for a concordant pair and ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=-1$ for a discordant pair, translating exactly the left-hand side into the definition of the Kendall. This means that changing the Kendall also implies that this is not a linear program anymore, thus MILP is a very rigid approach. Note that MILP is used as a standalone baseline without divide-and-conquer. In our implementation, MILP uses the COIN (CBC) solver (John Forrest et al., 2024).
Principal Component Analysis PCA
Principal Component Analysis (PCA) is a technique used to reduce dimensionality while minimizing loss in terms of variance. We use PCA on the performance matrix and look at the most important vector, the one who explains most of the variance. Then we choose to remove the test that is least important for this vector, that are the one whose coefficient is the least correlated with this vector.
5.1. Failed Attempts
In this section, we report other attempts to obtain some baselines that were not included in our experimental results. We decided not to include them due to poor results compared to existing baselines but believe it is of scientific interest to report what we failed to make work.
Racing Algorithm
Based on racing algorithms such as irace (López-Ibáñez et al., 2016), we adapt their approach based on Friedman’s test (and, 1937) and produce the same decisions irace would take. In other words, it works iteratively and chooses a test to remove at each step. The test chosen is based on Friedman’s test used in irace to detect the poorest performers; here, our intuition is that it should detect the test that has the least discriminative power and remove it. We implemented it in such a way that if we run irace on our setup, except that it does not have new data at each step, then irace would remove the same tests in the same order. Applying it in practice over all our benchmarks led to no cost reduction at all on all seeds; therefore, this baseline is discarded.
Genetic Algorithm
There is a large body of work on the minimization of the test suite that uses genetic algorithms (Harman and Jones, 2001; Yoo and Harman, 2012). However, most of these approaches are white- or gray-box approaches. Unlike these approaches, there are neither features nor associated code; therefore, we cannot adapt their approaches. We can still adapt the approach of genetic algorithms to our case. Our genes consist of zeros and ones; a one indicates that we keep a test in the test suite, and a zero indicates that it is discarded. Then, we can frame the RTSM problem as a multi-objective optimization problem by considering two independent objectives, namely 1) minimizing the size of the test suite and 2) minimizing the ranking error so that it stays below the desired threshold. Genetic algorithms, unlike other methods, cannot guarantee that the ranking error stays below the desired threshold; that is why we have to consider it as an objective function to minimize it. In order for genetic algorithms to work with our two objectives, we use the following formula:
$score(T)=\begin{cases}accuracy(T)\ \text{if $T$ is not accurate}\\ 10+(1-\frac{size(T)}{N})\end{cases}$
When we say ‘$T$ is not accurate’, we mean that it does not predict the classification of the variants to the required accuracy, that is, there is too much classification error. The idea is that we favor the sets $T$ that match the required accuracy threshold and then rank them according to their size, since $\frac{size(T)}{N}$ describes the ratio of tests kept from the problem. It is close to a linear combination, but since we want a solution in a limited time, we observed that we need to bias for correct solutions with the desired accuracy, to always get solutions within the timeouts.
We tried changing the score to a linear combination or using a multi-objective genetic algorithm, and we tried different variations of the parameters of the genetic algorithm to no avail: we only found worse-performing versions. Since we did not make genetic algorithms work correctly, they are not included in the experiments.
5.2. Practical Details
6. Experiments
We ask the following questions:
  • RQ1: How many tests can be eliminated while maintaining the same ranking?
  • RQ2: How many tests can be eliminated while maintaining approximately the same ranking?
  • RQ3: How is minimization affected with only a subset of the variants?
  • RQ4: How does our method compare to the others?
  • RQ5: How does the different components affect performance?
6.1. Experimental details
We implemented our methods in Python; an archive of the code is available as supplementary material along with all additional figures and tables. We will report Kendall’s $\tau$ as mentioned previously (see section 3). All experiments were run on a single core on Intel Xeon Gold 6130 (16 cores, 2017), and our algorithms returned the best result found so far. Working with multiple cores can speed up the results almost linearly, and better results may be achieved with a longer timeout on a single core, since all methods except MILP are anytime algorithms. Since our methods are stochastic, we conducted our experiments on 10 different seeds and reported mean performance along with the confidence interval 95% and performed statistical tests when necessary. The timeouts were 50 minutes per task for all datasets. As a reminder, all methods use divide-and-conquer and iterative solving except MILP.
6.2. Datasets Used
In general, we evaluate our methods on more than 50 benchmarks across a wide range of domains and software.
Code Generation: HumanEval (Chen et al., 2021; Liu et al., 2023b)
First, we looked at the HumanEval benchmark (Chen et al., 2021) and also considered the HumanEval Plus (Liu et al., 2023b) version with additional tests to correctly test generated code; furthermore, in the repository, the authors provide the samples generated by numerous models. We took these models as variants and the HumanEval tasks as tests, which gives us a matrix of zeros and ones depending on the pass $k$ score used. We took pass1 for both the classic version and the plus versions, giving us 2 benchmarks. These datasets contain 120 variants and 164 instances. The cost for these benchmarks is 1 for all tests, as no cost was provided in the original data.
Algorithm Selection: ASlib (Bischl et al., 2016)
Second, we looked at comparisons of algorithm performance. There are a large number of scenarios in the ASlib benchmark (Bischl et al., 2016), and we selected only scenarios where there were at least 8 variants. In fact, when too few variants are available, it is not as relevant. When multiple versions of the same data are available, we only keep the one with the most data. This produced 23 benchmarks with one metric to optimize, and the mean running time of the test was used as a cost; this means that these are instances of the WRTSM problem. The idea of using cost is to use the fact that some more complicated instances can provide more information but at the cost of more runtime than, for example, 10 smaller tests; therefore, not adding weights would be an unfair comparison of tests. This leads to a range of diverse benchmarks, from answer set programming, constraint solving, SAT, and others, with fewer than 100 variants and fewer than 500 tests. The number of variants is usually in the range of 8-20. The cost in these benchmarks is the mean running time of the tests in all variants.
Configurable Systems (Lesoil et al., 2024; Alves Pereira et al., 2021; Kaltenecker et al., 2020; Lesoil et al., 2023)
Third, we looked at datasets that measure the performance of configurable systems (Lesoil et al., 2024; Alves Pereira et al., 2021; Kaltenecker et al., 2020; Lesoil et al., 2023). Specifically, the dataset contains 8 unique benchmarks, but multiple metrics are measured; therefore, we generated variants of the benchmarks with an increasing number of metrics, but we limited ourselves to 4 variants of the same benchmark. This adds a pool of 19 additional benchmarks with up to 200 variants and thousands of tests. For sqlite we made only 4 versions despite 15 different output metrics. The cost in these benchmarks is the mean running time of the tests across all variants when available; if time is a metric, then it is also used as cost. For example, gcc-ctime has ctime as a cost and metric, whereas lingeling-conflicts has cost 1 for all tests since no time is in the metrics.
Program Repair: RepairBench (Silva and Monperrus, 2024)
Finally, we looked at the performance of LLM in automatic program repair with RepairBench (Silva and Monperrus, 2024), which contained two benchmarks at the time: gitbugjava (90 tests, 23 variants) and defects4j (484 tests, 22 variants). A variant here is an LLM. Then, three versions of each benchmark were created based on the release dates of the LLMs. The first cut-off date is Oct. 2024 (9 variants left), and the second one is Jan. 2025 (17 variants left). Adding a final six additional benchmarks. The cost for these benchmarks is 1 for all tests, since the cost was global and not on a test basis.
6.3. RQ1: How many tests can be eliminated while maintaining the same ranking?
In order to tackle this question, we will run our method on all datasets targeting no ranking error. The results are in LABEL:table:gain. For 17 of the benchmarks, no cost reduction was found, which means that for 33/50 we found a cost reduction. Finding no cost reduction is interesting because it means that all test instances are relevant for the ranking of the variants. From another point of view, there are a number of benchmarks where the cost reduction reaches near 1, so that means that almost all tests are useless to rank the variants; only a few are relevant. Looking more specifically at the data, BISS generally outperforms other approaches, notable contenders are RANDOM and MILP. This is why in future experiments we only consider MILP, BISS and RANDOM. It may surprise the reader that MILP is not the best performing solution since, by definition, it finds an optimal solution; this is because we cannot encode the full problem in MILP as such MILP solves a more constrained problem. As was explained previously, MILP indeed finds an optimal solution to the more constrained problem, which is not always the best solution to the true problem.
Table 1. Mean Cost reduction of different methods with all variants with 95% confidence interval in parenthesis if greater than 0. The target Kendall is 1. Higher is better. Best performing methods are in bold.
Fraction MILP BISS PCA RANDOM GREEDY
ASP-POTASSCO 0.00 0.99 (0.02) 0.27 (0.08) 0.80 (0.27) 0.23 (0.08)
BNSL-2016 1.00 1.00 0.53 (0.03) 1.00 0.53 (0.05)
CSP-Minizinc-Obj-2016 0.94 0.61 (0.27) 0.11 (0.13) 0.22 (0.07) 0.12 (0.06)
CSP-Minizinc-Time-2016 0.96 0.75 (0.27) 0.11 (0.08) 0.27 (0.08) 0.13 (0.08)
CSP-MZN-2013 0.00 1.00 0.19 (0.21) 0.94 (0.33) 0.11 (0.11)
defects4j 0.95 0.62 (0.52) 0.00 0.10 (0.04) 0.00
defects4j-at-2024-10-01 0.98 0.92 (0.07) 0.16 (0.04) 0.99 0.09 (0.05)
defects4j-at-2025-01-01 0.96 0.58 (0.52) 0.02 (0.04) 0.15 (0.04) 0.01 (0.03)
gcc-ctime 0.00 0.00 0.00 0.00 0.00
gitbugjava 0.00 0.00 (0.01) 0.00 0.00 0.00
gitbugjava-at-2024-10-01 0.86 0.35 (0.14) 0.01 (0.02) 0.23 (0.15) 0.00 (0.01)
gitbugjava-at-2025-01-01 0.00 0.08 (0.09) 0.00 0.01 (0.02) 0.00 (0.01)
GLUHACK-2018 0.00 0.97 (0.07) 0.35 (0.08) 0.98 (0.08) 0.21 (0.04)
humaneval-pass1 0.00 0.00 0.00 0.00 0.00
humaneval-plus-pass1 0.00 0.00 0.00 0.00 0.00
imagemagick-time 0.00 0.03 (0.01) 0.00 0.07 (0.07) 0.00
imagemagick-time-size 0.00 0.00 0.00 0.00 0.00
IPC2018 0.99 0.90 (0.18) 0.26 (0.09) 0.64 (0.15) 0.18 (0.08)
lingeling-conflicts 0.00 0.20 (0.02) 0.09 (0.05) 0.10 (0.04) 0.05 (0.04)
lingeling-conflicts-reduct. 0.00 0.00 0.00 0.00 0.00
MAXSAT-PMS-2016 1.00 0.96 (0.07) 0.36 (0.07) 0.72 (0.20) 0.29 (0.18)
MAXSAT-WPMS-2016 1.00 0.97 (0.04) 0.34 (0.19) 0.97 (0.05) 0.26 (0.08)
MAXSAT15-PMS-INDU 0.00 0.77 (0.23) 0.17 (0.16) 0.47 (0.05) 0.13 (0.15)
nodejs-ops 0.00 0.77 (0.32) 0.11 (0.18) 0.33 (0.19) 0.06 (0.03)
OPENML-WEKA-2017 0.00 0.59 (0.47) 0.00 0.17 (0.03) 0.00
poppler-time 0.00 1.00 0.37 (0.31) 0.93 (0.21) 0.38 (0.28)
poppler-time-size 0.00 0.00 0.00 0.00 0.00
PROTEUS-2014 0.00 0.55 (0.07) 0.08 (0.11) 0.27 (0.14) 0.07 (0.09)
QBF-2014 1.00 1.00 0.18 (0.15) 0.93 (0.38) 0.17 (0.11)
QBF-2016 0.00 0.63 (0.45) 0.12 (0.14) 0.31 (0.07) 0.08 (0.11)
SAT03-16-INDU 0.00 1.00 0.40 (0.28) 0.94 (0.14) 0.19 (0.15)
SAT11-INDU 1.00 0.75 (0.37) 0.20 (0.09) 0.36 (0.06) 0.14 (0.05)
SAT12-ALL 0.00 0.24 (0.09) 0.02 (0.02) 0.17 (0.14) 0.02 (0.06)
SAT15-INDU 0.00 0.40 (0.24) 0.10 (0.10) 0.21 (0.05) 0.09 (0.05)
SAT16-MAIN 0.97 0.76 (0.11) 0.25 (0.07) 0.35 0.16 (0.07)
SAT18-EXP 0.00 0.70 (0.28) 0.10 (0.11) 0.24 (0.05) 0.14 (0.07)
SAT20-MAIN 0.00 0.44 (0.15) 0.11 (0.08) 0.21 (0.08) 0.09 (0.07)
sqlite-q1 0.00 0.17 (0.16) 0.03 (0.06) 0.07 0.02 (0.06)
sqlite-q1-q5 0.00 0.00 0.00 0.00 0.00
sqlite-q1-q5-q10 0.00 0.00 0.00 0.00 0.00
sqlite-q1-q5-q10-q15 0.00 0.00 0.00 0.00 0.00
TTP-2016 0.00 0.84 (0.20) 0.05 (0.10) 0.31 (0.30) 0.03 (0.06)
x264-etime 0.00 0.00 0.00 0.00 0.00
x264-etime-cpu 0.00 0.00 0.00 0.00 0.00
x264-etime-cpu-size 0.00 0.00 0.00 0.00 0.00
x264-etime-cpu-size-fps 0.00 0.00 0.00 0.00 0.00
x264-etime-cpu-size-fps-kbs 0.00 0.00 0.00 0.00 0.00
xz-time 0.41 0.06 (0.13) 0.00 0.28 (0.29) 0.00
xz-time-size 0.00 0.00 0.00 0.00 0.00
Mean 0.27 (0.83) 0.44 (0.80) 0.10 (0.28) 0.30 (0.69) 0.08 (0.23)
Mean on reduced scenarios 0.38 (0.91) 0.63 (0.67) 0.15 (0.29) 0.43 (0.68) 0.12 (0.25)
In more than 66% of the benchmarks, test instances were successfully removed with high Kendall. Reductions of costs range from a few percent up to almost 100% of the cost, with the best method averaging 44% cost reduction on all benchmarks and 66% on benchmarks where a reduction was found by any of the methods. MILP, BISS and RANDOM largely outperform other methods.
6.4. RQ2: How many tests can be eliminated while maintaining approximately the same ranking?
For this RQ, we follow the same procedure as for RQ1, but with a Kendall target of 0.99, which allows for some leeway into the ranking while still being close to the original ranking. MILP is not present because as explained because we could not adapt the formulation to Kendall different from 1 without compromising the linearity.
The results are in LABEL:table:gain_99. For 13, down from 17, of the benchmarks, no cost reduction was found, which means that for 37/50 we found a cost reduction. We observe the same trends as for LABEL:table:gain. For PCA there is no improvement; instead, the mean reduction decreased. For RANDOM there is also no improvement and only a decrease in variance. For our method BISS, there is a 10% improvement over the mean.
Table 2. Mean Cost reduction of different methods with all variants with 95% confidence interval in parenthesis if greater than 0. The target Kendall is 0.99. Higher is better. Best performing methods are in bold.
Fraction BISS PCA RANDOM
ASP-POTASSCO 0.89 (0.30) 0.26 (0.10) 0.76 (0.31)
BNSL-2016 1.00 0.43 (0.31) 0.99 (0.02)
CSP-Minizinc-Obj-2016 0.67 (0.29) 0.15 (0.17) 0.23 (0.14)
CSP-Minizinc-Time-2016 0.63 (0.28) 0.09 (0.09) 0.24 (0.11)
CSP-MZN-2013 0.76 (0.71) 0.14 (0.08) 0.69 (0.50)
defects4j 0.70 (0.57) 0.01 (0.02) 0.16 (0.19)
defects4j-cutoff-2024-10-01 0.74 (0.21) 0.09 (0.13) 0.51 (0.77)
defects4j-cutoff-2025-01-01 0.51 (0.43) 0.02 (0.04) 0.24 (0.26)
gcc-ctime 0.12 (0.02) 0.00 0.14 (0.07)
gitbugjava 0.08 (0.13) 0.00 0.02 (0.05)
gitbugjava-cutoff-2024-10-01 0.20 (0.25) 0.00 (0.01) 0.17 (0.21)
gitbugjava-cutoff-2025-01-01 0.14 (0.20) 0.00 0.02 (0.04)
GLUHACK-2018 0.91 (0.20) 0.21 (0.24) 0.87 (0.41)
humaneval-pass1 0.00 0.00 0.00
humaneval-plus-pass1 0.00 0.00 0.00
imagemagick-time 0.97 (0.01) 0.05 (0.06) 0.73 (0.21)
imagemagick-time-size 0.00 0.00 0.00
IPC2018 0.82 (0.25) 0.20 (0.15) 0.46 (0.27)
lingeling-conflicts 0.29 (0.06) 0.07 (0.04) 0.09 (0.06)
lingeling-conflicts-reductions 0.00 0.00 0.00
MAXSAT-PMS-2016 0.91 (0.20) 0.22 (0.27) 0.65 (0.52)
MAXSAT-WPMS-2016 0.91 (0.14) 0.30 (0.24) 0.93 (0.33)
MAXSAT15-PMS-INDU 0.97 (0.02) 0.19 (0.26) 0.58 (0.50)
nodejs-ops 0.89 (0.54) 0.09 (0.05) 0.45 (0.44)
OPENML-WEKA-2017 0.77 (0.15) 0.08 (0.11) 0.28 (0.18)
poppler-time 1.00 0.23 (0.29) 0.84 (0.35)
poppler-time-size 0.00 0.00 0.00
PROTEUS-2014 0.68 (0.23) 0.06 (0.04) 0.24 (0.17)
QBF-2014 0.98 (0.08) 0.12 (0.08) 0.60 (0.63)
QBF-2016 0.73 (0.42) 0.10 (0.07) 0.32 (0.20)
SAT03-16-INDU 0.98 (0.08) 0.25 (0.09) 0.74 (0.45)
SAT11-INDU 0.59 (0.23) 0.13 (0.13) 0.25 (0.15)
SAT12-ALL 0.63 (0.40) 0.04 (0.03) 0.22 (0.20)
SAT15-INDU 0.45 (0.27) 0.10 (0.07) 0.18 (0.10)
SAT16-MAIN 0.75 (0.14) 0.19 (0.15) 0.32 (0.12)
SAT18-EXP 0.70 (0.32) 0.08 (0.06) 0.24 (0.11)
SAT20-MAIN 0.66 (0.26) 0.07 (0.06) 0.21 (0.10)
sqlite-q1 0.35 (0.10) 0.09 (0.06) 0.14 (0.09)
sqlite-q1-q5 0.00 0.00 0.00
sqlite-q1-q5-q10 0.00 0.00 0.00
sqlite-q1-q5-q10-q15 0.00 0.00 0.00
TTP-2016 1.00 0.08 (0.05) 0.77 (0.50)
x264-etime 0.68 (0.54) 0.04 (0.03) 0.46 (0.31)
x264-etime-cpu 0.00 0.00 0.00
x264-etime-cpu-size 0.00 0.00 0.00
x264-etime-cpu-size-fps 0.00 0.00 0.00
x264-etime-cpu-size-fps-kbs 0.00 0.00 0.00
xz-time 0.17 (0.18) 0.00 0.23 (0.23)
xz-time-size 0.00 0.00 0.00
Mean 0.49 (0.78) 0.09 (0.22) 0.31 (0.64)
Mean on reduced benchmarks 0.67 (0.61) 0.12 (0.23) 0.42 (0.62)
In more than 74% of the benchmarks, test instances were successfully removed with high Kendall. There is no improvement for PCA and RANDOM whereas losing .01 of Kendall increases the mean performance by 10% for BISS indicating that slightly decreasing Kendal provides huge benefits.
6.5. RQ3: How is minimization affected with only a subset of the variants?
For this RQ, as mentioned above, we only consider MILP, BISS and RANDOM. We sampled a fraction of all the variants for each benchmark, then ran the minimization procedure, and measured the cost reduction and Kendall on the full set of variants. Unlike RQ2, we only consider targeting a Kendall of 1, the same ranking, because we are already losing information by removing variants. This experiment reproduces a natural competition using our approach. The benchmark is minimized with only a subset of all variants, then new variants are submitted by competitors, and the reduced benchmark is used. We chose to keep 75% 50%, and 25% of all variants. We sampled 10 random subsets for each percentage and report the mean performance with a confidence interval 95%.
However, there are two metrics to consider: the cost reduction and the Kendall. Intuitively, if the cost is reduced by 10%, for an optimal trade-off, the Kendall should not be more than 10% lower. Therefore, we introduce a single metric that mixes the two in order to easily compare and run statistical tests on the performance of our methods. This metric is the score and gives equal weight to a cost reduction and a Kendall reduction: $score(c,k)=\frac{2c+(1+k)}{4}$ where $c\in[0;1]$ is the fraction of cost reduced and $k\in[-1;1]$ is the Kendall coefficient. By default, if no reduction is made, the obtained score is 0.5; another way to think about this is that if the score is above 0.5, then the trade-off starts to become interesting.
Since there are 50 different benchmarks, we produced 4 plots and a table for each benchmark, which are available in the supplementary material, since these would take more than 50 pages of PDF. As a general trend, the lower the fraction of variants kept, the greater the variance observed among the results. Surprisingly, we could expect the score to decrease with the lower fractions, but that is not always the case. Generally, the MILP approach has a lower variance than our approaches.
Instead, we look at the summary offered by Fig. 3, where we plot the cumulative distribution of the scores between seeds and benchmarks. The goal is to have a higher score, so the smaller the area under the curve, the better the method. Also note that on average, if a proportion greater than 0.5 is reached before a score of 0.5, then on average it means that the method is not worth it: we lose more than we gain either in terms of Kendall or in terms of reduction. Interestingly, this is the case for MILP, which is not worth it at all for different fractions of variants. In contrast, BISS and RANDOM are below this threshold, and are generally worth it.
In all scenarios, it is clear that the area under the curve is minimal for BISS. An interesting phenomenon occurs: the area under the curve tends to increase as the fraction of variants increases. This may seem counterintuitive at first; however, what happens is that removing variants often enables our approaches to drastically reduce the cost while losing little to no Kendall, therefore increasing the score. In other words, it is often the case that a benchmark cannot be reduced further because of a few variants that are hard to rank.
More generally, BISS offer scores of above:
  • 0.9 in more than 10% of cases for 25% of variants, 20% of cases for 50% of variants and 25% of cases for 75% and 100% of variants;
  • 0.8 in more than 40% of cases for all fractions of variants;
  • 0.5 in more than 65% of cases for all fractions of variants.
The last item is not surprising indeed; with 100% of variants, only 66% of the benchmarks could be reduced. Finally, in some scenarios, better solutions can be found when there are fewer variants than when given all variants, in other words, sometimes with less data, sometimes better solutions can be found.
(a) 25% of variants
(b) 50% of variants
(c) 75% of variants
(d) 100% of variants
Figure 3. Cumulative distribution of scores on the 50 benchmarks for different fraction of variants. Smaller Area Under the Curve is better.
BISS and RANDOM generally offer interesting trade-offs between cost and Kendall, with a significant portion of them having a high score (0.8 or above), making them quite worthwhile. MILP, on average, does not offer interesting trade-offs since more than 50% of the solutions have a score less than or equal to 0.5.
6.6. RQ4: How does our method compare to the others?
Using data from RQ1 and RQ3, we performed statistical tests and measured the effect size when comparing the scores of our method with those of all other methods. We ran a one-sided Wilcoxon signed-rank test to compare the scores of the different methods and also report the rank biserial correlation coefficient, that is, the effect size:
  • BISS outperforms RANDOM with effect size of 0.52 and p-value ¡ $10^{-277}$
  • BISS outperforms MILP with effect size of 0.24 and p-value ¡ $10^{-308}$
  • RANDOM outperforms MILP with effect size of 0.22 and p-value ¡ $10^{-173}$
BISS outperforms all other methods, RANDOM outperforms MILP, and then MILP outperforms the other methods.
BISS outperforms all other methods; however, we note on a few benchmarks that using other approaches may lead to better solutions; more precisely, the MILP approach offers new perspectives for some benchmarks, whereas RANDOM fails on the same benchmarks as our method.
6.7. RQ5: How does the different components affect performance?
Iterative solving
In order to measure the impact of iterative solving, we counted each time iterative solving improved the final solution over all our benchmarks for BISS with a Kendall target of 1 with 100% of variants, so a subset of the data from RQ1. We plot the cumulative distribution of the number of iterations performed on Fig. 4. We found that for most benchmarks iterative solving enables us to find better solutions in 76% of the cases where a cost reduction is found. Iterative solving is useless when the best solution is found during the first iteration, usually when a scenario is either very easy with only a few tests remaining or when a scenario is very hard with little to no tests being removed.
Figure 4. Cumulative distribution of iterations for BISS with a target Kendall of 1 with 100% of variants on the $3$ benchmarks where a cost reduction was found. Smaller Area Under the Curve is better.
Timeouts with Divide and Conquer
As part of our experiments, we recorded current solutions even in case of timeouts; when a timeout occurs, we use the full set of tests as a solution.
In order to measure, the impact of divide and conquer on the speed of our method, for all methods with divide and conquer on all different experiments we recorded the total number of timeouts. Then we looked for the number of timeouts of BISS and RANDOM without divide and conquer and iterative solving on the same scenario as RQ1, that is with all variants and with a Kendall of 1, this constitutes 1099 runs compared to the 58445 that we have with divide and conquer. As seen in Table 3, there is a large variation in the number of timeouts. Note that since MILP is not seed dependent, fewer runs are explaining the difference compared to the number from raw number to percentage.
First, not using divide and conquer multiplies by 2 the number of timeouts in number while having x50 time less runs, so not using divide and conquer multiplies by x100 the chance of timeout. Second, the best performing methods, BISS and RANDOM, have few timeouts (¡1%), MILP timeouts rarely. PCA surprisingly times out quite often.
Method BISS RANDOM PCA GREEDY MILP
Raw Number 144 40 5093 0 2650
Percentage 0.25% 0.07% 8.73% 0.00% 4.54%
Raw Number (No DC) 295 288 / / /
Percentage (No DC) 26.84% 26.21% / / /
Table 3. Timeouts per method used over all different seeds. No DC stands for the same results without divide and conquer and iterative solving. A ”/” stands for no data recorded. Note that only one experiment was run with No DC hence the different percentages. Since MILP does not depend on a seed there are also less runs for MILP.
Iterative solving improved the solutions in 76% of the benchmarks where we found a cost reduction. Divide and conquer reduces chances of timeouts by a factor of 100. Overall, for the best performing methods, BISS and RANDOM, there are very few timeouts with divide and conquer (¡1%) whereas there is one in four run that timeouts without it.
7. Selected Case Studies
In this section, we will look deeper into what we can learn from our solution:
  1. (1)
    BNSL: Is the dataset highly redundant?
  2. (2)
    SAT 18 and SAT 20: How does the test feature distribution change after minimization?
  3. (3)
    defects4j: Can benchmark minimization tell us about the evolution of LLMs’ skill on program repair?
7.1. Case Study: BNSL: Is the dataset highly redundant?
In (Matricon et al., 2021), the authors find that the BNSL benchmark from (Bischl et al., 2016) is easy to predict and emit the hypothesis that the benchmark is highly redundant. We want to confirm this hypothesis; therefore, we will look more precisely at the BNSL data extracted from our methods BISS and RANDOM the top performing ones on this scenario.
In Table 4, we give the mean performance and the standard deviation of the different score for the different methods, and on Fig. 5 we instead plot the different data points that we have, that is we plot the cost ratio with respect to the Kendall obtained for different fractions of variants. First, observe in Fig. 5 that all our methods cluster around a very small cost ratio; only a few data points are above a cost ratio of 0.001, leading us to think the test instances are highly redundant. Except for MILP, on Table 4, all scores are very high, however there is a high variance when variants are removed, indeed since the cost reduction is so high: it is because only one test is kept, however the selected test highly depends on the subset of variants kept thus increasing the variance. Overall, looking at the data, their benchmark is highly redundant as was hypothesized in (Matricon et al., 2021).
Fraction BISS RANDOM
100 0.99 (0.05) 0.93 (0.11)
75 0.92 (0.14) 0.88 (0.11)
50 0.88 (0.17) 0.83 (0.17)
25 0.88 (0.17) 0.81 (0.20)
Table 4. Scores with 95% confidence interval in parentheses of different methods for BNSL-2016. Best performing methods are in bold.
(a) 2 (25%) variants
(b) 4 (50%) variants
(c) 6 (75%) variants
(d) 8 (100%) variants
Figure 5. Kendall (higher is better) with respect to cost ratio (lower is better) for different methods across different fractions of variants kept for BNSL-2016 with a target Kendall of 1
BNSL, as hypothesized in (Matricon et al., 2021), is a highly redundant dataset with a ranking of variants that can be done with only one test instance.
7.2. Case Study: SAT 18 and SAT 20: How does test features distribution change after minimization?
In this case study, we want to observe how the different statistical features of the tests evolve when we only select a subset of the tests. We will use the 10 different solutions with 100% of variants obtained with BISS. The minimization likely induces a significant change in the test set. We use the features given in (Bischl et al., 2016) that were used for automated algorithm selection tasks. Intuitively, if the variants are representative and the features relevant, then we should keep more or less similar distributional properties on the reduced test sets. To check this, we look at the evolution of min, 1st quartile, median, 3rd quartile, max, mean, and std for all features. We compute the median relative evolution of these characteristics across all features. The median has the interesting property that if only a few features vary wildly, this will not be seen in the median, and at the same time, if globally there are a lot of variations, the median will capture that. We selected two datasets with varying minimization successes despite being from the same domain. For SAT18, there are 54 features; we found cost reductions ranging from 45% up to 85%, which translates into keeping only between 38% and 85% of the original tests. For SAT20, there are 108 features; we found cost reductions ranging from 38% up to more than 60% which translates into keeping only between 24% and 72% of the original tests.
The results are in Table 5 for SAT 18 and Table 6 for SAT 20. Interestingly, in both cases, there almost always seems to be no change to the minimum,m indicating that perhaps it is almost always a good choice to keep it. We could expect the same for max, but it is not as clear, it still seems to be very relevant since there are a lot of 0% but that is not always the case. More precisely, the change from the 1st quartile to the median to the 3rd quartile is always increasing. Intuitively, larger values for features usually translate into more complex instances, which are slower to solve; therefore, removing them provides greater cost reductions so it is likely that our approach focuses on removing these instances first. The best minimization seems to have no relative changes in min or max.
For SAT 20, it is easy to observe that there are very few changes to the distributional property, with the highest relative changes never going over 9%. For SAT 18, despite smaller changes to the test set than for SAT 20, there are more impactful changes up to 23% except for the 3rd quartile and max with changes of 36.87% and 26.43% respectively, this is likely due to greater cost reduction obtained for this benchmark.
percentage of tests kept mean std min Q1 median Q3 max
44.48% 8.79% 4.94% 0.00% 3.90% 7.02% 18.86% 0.00%
72.52% 11.83% 8.16% 0.00% 3.72% 9.98% 36.87% 4.40%
74.79% 9.04% 10.59% 0.00% 5.47% 8.32% 17.22% 3.09%
49.29% 8.62% 7.07% 0.00% 8.50% 8.33% 28.17% 0.00%
84.99% 5.18% 9.04% 8.18% 4.76% 5.77% 7.03% 26.43%
76.77% 7.79% 8.67% 0.00% 5.39% 10.33% 13.64% 10.70%
78.19% 6.09% 5.99% 0.00% 6.01% 6.35% 11.39% 10.22%
82.15% 12.98% 8.35% 0.00% 3.68% 7.67% 15.09% 9.51%
38.81% 4.38% 4.26% 0.00% 4.68% 5.84% 17.83% 0.00%
80.17% 9.16% 6.29% 0.00% 3.25% 7.52% 12.70% 25.78%
Table 5. Relative median change across features with one seed per row for solutions of BISS on SAT 18. For each column the largest value is in bold and the smallest is underlined.
percentage of tests kept mean std min Q1 median Q3 max
39.06% 4.43% 4.50% 0.00% 2.91% 2.01% 3.90% 0.00%
24.41% 3.62% 4.30% 0.00% 1.63% 1.55% 2.17% 0.00%
71.38% 8.60% 8.92% 0.00% 3.58% 7.98% 7.92% 4.17%
27.65% 4.30% 5.71% 0.00% 1.39% 1.90% 2.69% 0.00%
24.58% 3.68% 3.55% 0.00% 0.67% 1.17% 2.50% 0.00%
36.70% 4.30% 5.36% 0.00% 1.85% 2.28% 3.46% 0.00%
25.59% 3.94% 3.45% 0.00% 0.93% 1.43% 1.60% 0.00%
65.73% 8.28% 7.62% 0.00% 3.05% 5.53% 8.81% 6.97%
33.82% 3.70% 3.95% 0.00% 1.33% 1.96% 2.75% 0.00%
26.18% 3.68% 3.13% 0.00% 1.31% 1.21% 3.32% 0.00%
Table 6. Relative median change across features with one seed per row for solutions of BISS on SAT 20. For each column the largest value is in bold and the smallest is underlined.
BISS keep distributional properties of most features of the tests, with most relative changes largely smaller than the ratio of tests kept.
7.3. Case Study: Defects4j: Can benchmark minimization tell us about the evolution of LLMs’ skill on program repair?
Defects4J is a benchmark of (Silva and Monperrus, 2024) and measures the abilities of LLMs to do program repair. To evaluate the evolution of the skills of LLMs, here, the variants, we plot on Fig. 6 the different data points of Kendall with respect to the cost ratio for our methods. The Kendall is measured on the full dataset, that is, with all LLMs. Here, the cost ratio is the ratio of tests kept in the minimized benchmark.
Interestingly, for the cutoff date in 2024, we observe a very low cost ratio, indicating that most LLMs before this date can be easily ranked. The Kendall, however, is quite low, indicating that the behavior of the more recent LLMs cannot be correctly captured by the present minimized test set. For the cutoff date in 2025, we observe that the Kendall jumps significantly, translating the fact there has been a lot of progress of LLMs on the task with for example Claude Sonnet 3.5 which is still one of the best performing models tested with a 33% relative increase in performance from the previous best, and that they behave more consistently with more recent LLMs. Note that at that date, there were no ‘reasoning’ LLMs; this is why we also observe that the Kendall si not yet close to 1, translating the new behavior captured by those new models. In terms of cost ratio, we observe heterogeneous results that are actually quite similar to the cost ratio distribution of the case with no cutoffs, which translates to the correct number of selected tests, but they are not the most relevant for ranking all LLMs. Finally, with no cutoffs, we observe that we can have cost reductions of more than 90% of the tests set with no loss of accuracy in the best case, indicating that the benchmark is quite redundant for measuring the performances difference among the LLMs considered. With MILP, we obtain a cost ratio of 5%, meaning that all the LLMs’ performances can be correctly ranked with only 5% of the tests, which can save hundreds of dollars per attempt according to the costs reported by (Silva and Monperrus, 2024).
(a) cutoff on 2024/10/01 (9 variants/LLMs)
(b) cutoff on 2025/01/01 (17 variants/LLMs)
(c) no cutoff (22 variants/LLMs)
Figure 6. Defects4J Kendall (higher is better) with respect to cost ratio (lower is better) with 100% of variants but with different cutoff dates with a target Kendall of 1
Defects4J contains a lot of redundant instances that bring little to no discriminative power to rank the different LLMs 5% of the 484 bugs are enough to perfectly rank all the LLMs, saving hundreds of dollars per attempt. The evolution of the cutoff dates and the results show that LLMs before Oct. 2024 behaved quite differently from the ones after, the LLMs released between Oct. 2024 and Jan. 2025 are more similar in terms of performance to the ones released afterwards.
8. Threats to Validity
Internal threats.
Timeouts
The configuration of timeouts presents a potential threat to the internal validity of our study, but since most methods are anytime, this leaves as much time as needed for the different methods.
Stability
While it is theoretically true that adding more tests to a solution can only improve it, in practice, this is not always the case. We have observed that adding more tests can sometimes lead to instability. For example, our $FindNecessary$ procedure (see Algorithm 1) may incorrectly identify certain elements as necessary when they are not. Unfortunately, we do not yet have a solution to this problem. However, our experimental results suggest that it is a minor issue. Our best experimental method is based on the $FindNecessary$ procedure, and the results show that the impact of this issue is very low.
Generalizability
The effectiveness of our method relies on several hyperparameters, particularly within the divide-and-conquer part. These hyperparameters include the number of splits used to manage variance and the number of tests conducted during the iterative solving phase. While these parameters were selected manually, we avoided overfitting and extensive hyperparameter tuning to maintain the generalizability and robustness of our method across different datasets and scenarios. By selecting reasonable default values, we aimed to balance performance and applicability. Besides, we repeated each experiment ten times to account for variability and random effects. This repetition did introduce variance in the results. We reported and analyzed the standard deviations across different runs as part of our findings.
External threats. The applicability of our study to real-world situations is influenced by the types of data we used for the evaluation. We tested our method using specific benchmarks from well-known sources, but this may not be representative of all possible datasets and domains. Our evaluation spans different domains (LLM, SAT, video, …), with variants being independent systems or configurations from the same system, and different scenarios (leaderboard, competitions, performance modeling). An empirical question remains regarding the extent to which benchmarks can be minimized in each domain while still maintaining accurate rankings. Addressing this question is an ambitious task that would require extensive experiments. Our paper highlights the need for further exploration and assessment across broader datasets. We hope that other researchers will build on our work and conduct additional studies.
9. Impact
In general, this analysis shows that some benchmarks are more amenable to minimization than others, highlighting the importance of domain specificities when using BISS.
Recommendations for benchmark design. Benchmark designers, developers, creators can use our methods to design, understand, and refine their benchmarks. Our method automates this process by retaining the most relevant ones and, as shown, effectively removing tests that are too easy or repetitive. It can be applied to existing benchmarks, effectively compressing them. We recommend using BISS since it generally outperforms other methods, and in case of bad results, to use MILP, which behaves quite differently. If benchmark designers want to test new variants on the reduced dataset (e.g., to verify the hardness of the benchmarks), they should replicate the RQ2 experiments to ensure representativeness and control ranking error. As experiments show, on average better results are achieved when most of the variants remain in the problem.
Recommendations for software developers Developers consume benchmarks for performance comparison, nonregression, or tuning. Smaller benchmarks with the same discriminative power mean that consumers will maintain accuracy while spending less time executing their software on variants. For scenarios like tuning or nonregression, where benchmarks are run repeatedly, the gains are amplified with each run, resulting in substantial overall benefits. We thus offer the same recommendations as for benchmark design, and empirical results show that important gains can be achieved for reducing the computational cost.
In addition, our method enables strategic targeting of the most performance-critical tests. While general recommendations remain, developers who need to focus on a very few instances can even reduce the target Kendall to find even smaller representative instances at the cost of accuracy. They can then improve their systems before testing on the larger benchmark. However, there is a risk of overfitting (e.g., BISS may find one test case sufficient to rank) and of excessively aggressive tuning or non-regression assurance. Developers should balance these risks and gradually increase their confidence based on the computational budget.
Cost Saving Potentials A highlight of our results is, for example, defects4j, which can be reduced to only 5% of the tests, saving up to hundreds of dollars per new LLM test. These reductions can be observed on many different datasets at different scales. Our results are available in the supplementary material, and since our tool is available, this enables benchmark creators to easily check their benchmark.
10. Conclusions and future work
We have introduced the ranked test suite minimization (RTSM) and its solution. It is a new problem that aims to minimize a benchmark while keeping the performance ranking of the variants over the test suites intact. This problem is highly relevant for practitioners, as it enables cost-effective benchmarking for variant comparison, rapid regression detection, performance modeling, benchmark design, etc.
Addressing this problem requires developing new approaches specifically designed to optimize benchmark selection while preserving variant ranking accuracy. We have introduced our approach BISS based on bisection sampling and supplemented with divide and conquer and iterative solving; and introduced baselines to compare to, including MILP, PCA, RANDOM, and others based on existing techniques. Experiments on 50 different benchmarks show that divide-and-conquer enables a 4-times faster solving and that iterative solving always leads to improved results. Iterative solving improves the solutions in 76% of the cases where a cost reduction was found. Our experiments show that BISS outperforms all alternatives. Furthermore, when reducing the target Kendall, we show that our approach highly benefits from it, by reducing the Kendall by 1% on average, we get a 10% improvement. We also showed that our method behaves well even if all variants are not present, it often introduces some imprecision; however, we showed that in most cases, this is an interesting trade-off, we gain more cost reduction than we lose Kendall or Spearman.
In general, with BISS, more than half of the benchmarks could be reduced with a mean reduction over these reducible benchmarks of 67%. This implies that quite extensive reductions are available for many benchmarks, including those used today. Future work includes studying the applications and implications of our methods in different domains. For example, the field of AI is driven by benchmark, standardizing such minimization approaches could drastically reduce the cost of research comparison, thus increasing research velocity. Another direction is to work with partial data, that is, with some variants running only on a subset of the tests.